OpenStudy (anonymous):

How do I determine the sum of this sequence: 2; 6; 18;...; 3,188,646?

OpenStudy (anonymous):

I set up the explicit formula as: 3188646=2*3^(n-1)

OpenStudy (anonymous):

then I subtracted 2 to get 318644=3\[^{n-1}\]

OpenStudy (anonymous):

\[3,188,644=3^{n-1}\]

OpenStudy (anonymous):

I next did log of both sides and divides both by log3, but the left side was 13.6 and some more decimals, where did I mess up?

OpenStudy (anonymous):

\[\frac{ \log_{3,188,644} }{ \log_{3} }=\frac{ (n-1) \log_{3} }{ \log_{3} }\]

OpenStudy (anonymous):

Now what?

OpenStudy (tkhunny):

#1 There is insufficient information. We have no idea how many terms or what sort of sequence it is. #2 If we are to ASSUME that it is a geometric sequence, THEN we can add it up. 2 = 2*3^0 6 = 2*3^1 18 = 2*3^2 ... 3,188,646 = 2*3^13 Okay, that's a geometric series with 14 terms. First Term: 2 Common Ratio: 3 Go!

OpenStudy (anonymous):

Yes, sorry, it is a geometric sequence. That way makes sense and thanks for showing it to me, but when I do the log way in my calculator I get: 13.63092918=n-1 Why isn't this a perfect 14?

OpenStudy (anonymous):

Okay, the 14=n makes sense, so I'll proceed with that. Now I do the geometric sum formula, no? S=\[S = \frac{ 2 (1-3^{14})}{ 1-3 }\]

OpenStudy (tkhunny):

The ONLY thing for which we might need the logarithm is to find the 13. You failed to remove the 2. \(2\cdot3^{n} = 3,188,646\) \(3^{n} = 1,594,323\) \(n\cdot\log(3) = \log(1,594,323)\) \(n = \dfrac{\log(1,594,323)}{\log(3)} = 6.2025763114/0.4771212547 = 13\)

OpenStudy (anonymous):

n-1=13, right?

OpenStudy (tkhunny):

There are 14 terms in the sequence. The greatest exponent in the sequence is 13. This makes 14 the correct value in that formula.

OpenStudy (anonymous):

\[S=\frac{ 2(1-4782969) }{ -2 }\] Next, \[S=1-4782969\] \[Final, S=4782969\]

OpenStudy (anonymous):

Thanks a ton, tk. You've helped me before, so I knew you would be helpful.

OpenStudy (anonymous):

errr, final =4782968

OpenStudy (tkhunny):

There are some sign problems in there, but you did manage to land in the right spot. Good work.